1. Given C=20+0.8YD, I=20+5i, G=500, X=400, and T=200.

a). Derive the IS equation

b). Find the value of the multiplier

c). Find the equilibrium consumption

2. Assuming Government spending increases by 50% and investment reduces by 10%, what happens to the IS equation



1
Expert's answer
2021-08-29T16:54:30-0400
  1. "C=20+0.8YD, I=20+5i, G=500, X=400, and T=200."

a.

IS equation is given by"y = c + i + g + x + t"

"y = 20 + 0.8yd + 20 +5i + 500 + 400"

but disposable income is"yd = y-t"

"y = 20 + 0.8(y-200) +20 + 5i +500+400"

Removing the brackets and Making y the subject

"0.2y = 5i +740"

hence "y= 25i + 3700"

b.

the value of the multiplier is the coefficient of i in the IS equation


multiplier = 25

c.

equilibrium consumption is given by

"c= 20 + 0.8yd"

"c = 20 + 0.8(y-200)"

but y =25i + 3700

Therefore "c= 20+ 0.8(25i + 3700 -200)"

"c = 20i + 2820"

2.

The proportion in the increase in government spending is substantially huge as compared to a decrease in the interest rate. Therefore, the net effect is that IS equation will broaden thereby shift outwards to right hand right and upwards.